OpenStudy (anonymous):

Carrie started with 12 ounces of water when she began her walk. when she finished walking she had 2 1/4 ounces if water. how much did she drink

OpenStudy (anonymous):

\[12-\frac{9}{4}=\frac{39}{4}=9.75=9+\frac{3}{4} \]

OpenStudy (mathstudent55):

This is a subtraction problem. Do you now how to subtract a mixed numeral from a whole number?

ganeshie8 (ganeshie8):

@dwilso22